OpenStudy (anonymous):

my answer: 2x2 - 4x - 3 - 4x - 3/13 ?????? question: Let f(x) = 8x3 - 22x2 - 4 and g(x) = 4x - 3. Find f(x)/g(x)

OpenStudy (anonymous):

am i right or am i right

OpenStudy (anonymous):

lol.

OpenStudy (anonymous):

@zepdrix, is she right? I think she may not be.

OpenStudy (anonymous):

FML

OpenStudy (anonymous):

well u have to try and simplify first so how did u factor the top?

OpenStudy (anonymous):

f(x)**

zepdrix (zepdrix):

So we can work this one out using Polynomial Long Division. I worked it out and I came up with, \[\large \frac{f(x)}{g(x)}=2x^2-4x-3, \qquad \text{with a Remainder of -13}\] Which we can write like this. \[\large \frac{f(x)}{g(x)}=2x^2-4x-3+\color{royalblue}{\frac{-13}{4x-3}}\] It looks like you had everything correct except for that last little part! :) Just remember to put your Remainder (left overs) on top of the initial Divisor that you had (4x-3 in this case).